Pikachu wants to open a store to sell pens and ramen. He must borrow $30,000 from Papa Smurf's crew at 105% interest per year. He must pay this money back to Papa Smurf in his first year. He also must rent a storefront for $25,000 per year and hire his friend Patrick Star and pay him $650,000. He quit his job which earned him $250,000 per year. In his first year, he sold 11,000 pens at $3 each and 2500 bowls of ramen at $6 each. A bowl of ramen costs $3 in raw material costs and each pen costs him $2 each. 


For his 1st year please provide (1) his total economic cost (2) his economic profit (3) his accounting profit (4) should he have quit his job?


1
Expert's answer
2021-03-23T07:30:04-0400

Solution:

1.). Total economic cost = Explicit costs + implicit costs (Opportunity costs)

Explicit costs = Loan interest + rent + wages + raw material costs

Loan interest = 30,000"\times" 105% = $31,500

Rent = $25,000

Wages = $650,000

Raw material costs = "(11000\times 2) +(2500\times 3)" = 22,000 + 7,500 = $29,500

Explicit costs = 31,500 + 25,000 + 650,000 + 29,500= $736,000

Implicit costs = opportunity costs foregone = $250,000

Total economic costs = 736,000 + 250,000 = $986,000

 

2.). Economic profit = Total revenue – explicit costs – implicit costs

Total revenue = "(11000\times 3) +(2500\times 6)" = 33,000 + 15,000 = $48,000

Economic profit = 48,000 – 736,000 – 250,000 = ($713,000)

Economic loss = ($713,000)

 

3.). Accounting profit = Total revenue – explicit costs

= 48,000 – 736,000 = ($688,000)

Accounting loss = ($688,000)


4.). Pikachu should not have quit his job. This is because of the huge losses that he made for the year, which includes accounting loss and economic loss, yet he could have earned $250,000 for the year.